# Using Contextual Information for Sentence-level Morpheme Segmentation

### v1

@@ 2.3 Training
Our model is trained using entmax-loss (Peters et al., 2019) with alpha of 1.5 similar to Peters and Martins (2022) since it seemed to show better performance compared to cross-entropy loss in our preliminary experiments. The basic idea behind entmax loss is to penalize the model when it produces probability distributions that are too peaked or overly concentrated on a single class. One advantage of entmax loss over cross entropy loss is that it encourages the model to produce probability distributions that are more evenly distributed over the possible classes. This can help the model make more nuanced predictions and avoid over-confident mistakes caused due to overly confident predictions on a single class. This makes entmax loss a better choice over cross-entropy loss in many applications. However, entmax loss is computationally more expensive to compute than cross-entropy loss. We therefore use existing implementation of entmax loss rather than implementing it from scratch.

### v2

@@ 2.3 Training
Our model is trained using entmax-loss (Peters et al., 2019) with alpha of 1.5 similar to Peters and Martins (2022) since it seemed to show better performance compared to cross-entropy loss in our preliminary experiments.
The entmax loss is a general family of loss functions encompassing the cross-entropy loss as a special case (when alpha=1). Using values greater than 1 for alpha, the entmax loss allows for sparse gradients with non-zero values occurring only on the gold label and/or other labels that receive non-zero probability. This means that completely-peaked probability distributions are possible, in contrast to the denser distributions observed at alpha=1.” This makes entmax loss a better choice over cross-entropy loss in many applications. However, entmax loss is computationally more expensive to compute than cross-entropy loss. We therefore use existing implementation of entmax loss rather than implementing it from scratch.